Marcel Co. is growing quickly. Dividends are expected to grow at a 18 percent rate for the next 3 years, with the growth rate falling off to a constant 4 percent thereafter.

  

Required:

If the required return is 8 percent and the company just paid a $1.90 dividend. what is the current share price? (Do not round your intermediate calculations.)
Answer #1

Dividend in year 1=(Dividend just paid)*(1+Growth rate)=1.9*(1+18%)=2.242
Dividend in year 2=1.9*(1+18%)^2=2.64556
Dividend in year 3=1.9*(1+18%)^3=3.1217608

Terminal value=[3.1217608*(1+Constant growth rate)]/(Required return-Constant growth rate)
=[3.1217608*(1+4%)]/(8%-4%)
=3.246631232/4%
=81.1657808

Current share price=(Dividend in year 1)/(1+required return)^1+(Dividend in year 2)/(1+required return)^2+(Dividend in year 3)/(1+required return)^3
+(Terminal value)/(1+required return)^3

Current share price=(2.242)/(1+8%)^1+(2.64556)/(1+8%)^2+(3.1217608)/(1+8%)^3+81.1657808/(1+8%)^3
=(2.242)/1.08+(2.64556)/1.1664+(3.1217608)/1.259712+81.1657808/1.259712
=2.075925926+2.268141289+2.478154372+64.43201367
=71.25423526 or $71.25 (Rounded to 2 decimal places)

Answer: Hence, the current share price is $71.25

Please let me know if you face any difficulties in understanding the solution.